OpenStudy (anonymous):

The equation 2x 2 −62x+k=0 has two real roots, one of which is 1 more than twice the other. Find the value of k .

OpenStudy (anonymous):

\[2x^{2}- 62x + k = 0\]

OpenStudy (anonymous):

help me

OpenStudy (loser66):

let m, n are roots,m -1=2n and you have formula m +n = -b/a and m*n = c/a replace and solve for m, n

OpenStudy (anonymous):

please what are the roots confusion

OpenStudy (anonymous):

soorry

OpenStudy (anonymous):

ohh wait help come backkkk

OpenStudy (anonymous):

@Loser66

OpenStudy (e.mccormick):

Do you mean what are roots, or what are the answers?

OpenStudy (anonymous):

no the roots not the answer i was wondering how is that related to the answer too I would never ask for theanswer just walk me through thats all

OpenStudy (anonymous):

if u give me the answer thats just like saying im dumb I hate it

OpenStudy (e.mccormick):

Well, the roots of an equation is another way of saying the zeros.

OpenStudy (anonymous):

Could just identify the roots in the problem plz

OpenStudy (anonymous):

:)

OpenStudy (e.mccormick):

In the problem, they are a range because of the k and the conditions. They are not individual values.

OpenStudy (e.mccormick):

Or... hmmm...

OpenStudy (e.mccormick):

"has two real roots, one of which is 1 more than twice the other. Find the value of k ." OK. Read that wrong at first... not a range. Those are two distinct values.

OpenStudy (anonymous):

All i though is the anwer is between the interger valus 0 to 900

OpenStudy (anonymous):

I mean 999

OpenStudy (anonymous):

hmm... this is super hard but I love it

OpenStudy (e.mccormick):

OK, so if you have these roots, \(r_0\) and \(r_1\), that thing is saying that: \(r_0=?\) and \(r_1=2(r_1)+1\) Now, because keeping track of sub this and sub that is a pain, and because +1 on one side is the same as -1 on the other, L66 wrote that as: Let m and n be the roots where: m -1=2n

OpenStudy (anonymous):

when u have (x-r1)(x-r2), they should give you the original equation

OpenStudy (e.mccormick):

Yes.

OpenStudy (anonymous):

okay so how do use this to get the answer

OpenStudy (anonymous):

WAIT, there's a far easier way to do this using the quadratic formula/discriminants

OpenStudy (e.mccormick):

Another thing L66 was pointing out was a relationship between the roots and the positions of things in the equation. Remember that these things are in the form of: \(ax^{2}+bx + c = 0\)

OpenStudy (anonymous):

I believe that using the quadratic formula would be far faster and easier to understand \[2(62-\sqrt{62*62-8k})/4+1=(62+\sqrt{62*62-8k})/4\]

OpenStudy (anonymous):

just solve for k in that equation

OpenStudy (anonymous):

oh im dumb actually

OpenStudy (anonymous):

okay hold wait sec plz im dumber than

OpenStudy (anonymous):

okay, in an quadratic where r and s are the roots, then k in the equation is just rs so far so good?

OpenStudy (anonymous):

keeping with the previous equation, (realized that im not that dumb lol) just solve for \[\sqrt{62*62-8k}\] , as a 1 unknown algebra equation

OpenStudy (anonymous):

32 - 8k + 1 = 32 - 8k

OpenStudy (anonymous):

oh wow after the trouble

OpenStudy (anonymous):

note that one is + and the other is -

OpenStudy (anonymous):

wait than wont k be 0

OpenStudy (e.mccormick):

Umm... I thinkl you just skipped the condition on the roots and their relationship.

OpenStudy (anonymous):

I'm getting k as 420

OpenStudy (anonymous):

he did

OpenStudy (anonymous):

let a be the sqrt(stuff) the u have 2*(62-a)/4 +1 = (62+a)/4

OpenStudy (anonymous):

i going check it its right than i goign write u a testmonial if its wrong all the credit goes to @e.mccormick

OpenStudy (e.mccormick):

He got the right answer. Hmm.

OpenStudy (anonymous):

theres also another k answer i believe

OpenStudy (anonymous):

no wait im dumb lol

OpenStudy (e.mccormick):

With those conditions, it would be hard to get 2. LOL

OpenStudy (anonymous):

I went in imaginary by accident lol

OpenStudy (anonymous):

so it 420 right

OpenStudy (e.mccormick):

Ah, but you mean if the \(\pm\) is reversed.

OpenStudy (anonymous):

im checking okay

OpenStudy (anonymous):

yah @e.mccormick

OpenStudy (anonymous):

but its just the negative of the answer, so the square is the same

OpenStudy (anonymous):

Alright 420 was right sorry i got these problems for brilliant

OpenStudy (anonymous):

wow that took more effort than expected lol

OpenStudy (anonymous):

Got brilliant it hard to get these answers

OpenStudy (e.mccormick):

Now, to use \(ax^{2}+bx + c = 0\) with \(2x^{2}- 62x + k = 0\) m*n=k/2 m+n=62/2 m = 2n+1 (2n+1)*n=k/2 2n+1+n=62/2 (2n+1)*n=k/2 3n=30 n=10 (2(10)+1)*10=k/2 (21)*10=k/2 210=k/2 420=k That was where I was going.....